OpenStudy (anonymous):

Terrel and Mindi are organizing a neighborhood skateboarding competition. They need to build two mini ramps. Terrell can build the ramps in 15 hours. Mindi can build the ramps in 20 hours. Which shows the equation and solution for the time it will it take Terrell and Mindi to build the ramps if they work together?

OpenStudy (nottim):

wat have u done so far

OpenStudy (anonymous):

@Calcmathlete

OpenStudy (anonymous):

i just am really confused

OpenStudy (nottim):

ok.

OpenStudy (anonymous):

\[\frac{1}{15} + \frac{1}{20} = \frac{1}{x}\]Solve for x.

OpenStudy (anonymous):

What I did is find how much work each does in one hour and set it equal to the amount of work done with the combined work rate in an hour.

OpenStudy (anonymous):

so 4 hours?

OpenStudy (anonymous):

It would be 60/7 hours which you can approximate it... \[\frac{4}{60} + \frac{3}{60} = \frac{1}{x}\]\[\frac{7}{60} = \frac{1}{x}\]\[7x = 60\]\[x = 60/7

OpenStudy (anonymous):



OpenStudy (anonymous):

Do you understand what I did?

OpenStudy (anonymous):

yep thanks much i have to round it to nine thanks

OpenStudy (anonymous):

np :)